media%2Ff07%2Ff07fcbe5-3f24-4752-a852-ef
Answer #1

7) 1 Kilogram = 1000000 milligrams

1 Kg = 1000 gm and 1 gm = 1000 mg, hence, 1 Kg = 1000 * 1000 mg = 1000000 mg

8) Density (d) = mass/volume

A. 139ml and 93 g \therefore d = 93/139 = 0.66907 g/ml

B. 13dm3 and 1290 g ; 1dm3 = 1000ml

\therefore d = 1290/13000 = 0.09923 g/ml

C. 2.5L and 12.5 kg; 2.5L = 2500ml; 12.5kg= 12500 g

\therefore d = 12500/2500 = 5 g/ml

D. 3.91*10-24 nm3 and 7.93*10-1 ng; 1 nm3 = 1*10-21 ml \therefore 3.91*10-24 nm3 = 3.91* 10-45 ml

and 1ng = 1*10-9 g; 7.93*10-1 ng = 7.93*10-10 g

\therefore d = 7.93*10-10 g/3.91* 10-45 ml = 2.02813*1035 g/ml

E. 1m3 = 1000000 ml \therefore 0.00212 m3 = 2120 ml

and 4.22*104 mg = 42.2 g

d= 42.2/ 2120 = 0.01991 g/ml

Hence the highest density is of D. (3.91*10-24 nm3 and 7.93*10-1 ng)

9) 1 picometer = 1*10-10 cm (Answer is E)

10) 1mg = 1000 micrograms

11) Density of TCE = 40.55/25 = 1.622 gm/ml and density of water is 1 gm/ml, hence, densuty of TCE more than water and hence, it will not float on water.

Density of SuperC. CO2 = 0.469 gm/ml, we know, d = mass/volume \therefore mass = 0.469*25 = 11.725 gm.